### v1

# Rethinking Backdoor Detection Evaluation for Language Models

Abstract
Backdoor attacks, in which a model behaves maliciously when given an attacker-specified trigger, pose a major security risk for practitioners who depend on publicly released language models. Backdoor detection methods aim to detect whether a released model contains a backdoor, so that practitioners can avoid such vulnerabilities. While existing backdoor detection methods have high accuracy in detecting backdoored models on standard benchmarks, it is unclear whether they can robustly identify backdoors in the wild. In this paper, we examine the robustness of backdoor detectors by manipulating different factors during backdoor planting. We find that the success of existing methods highly depends on how intensely the model is trained on poisoned data during backdoor planting. Specifically, backdoors planted with either more aggressive or more conservative training are significantly more difficult to detect than the default ones. Our results highlight a lack of robustness of existing backdoor detectors and the limitations in current benchmark construction.
Rethinking Backdoor Detection Evaluation for Language Models
Jun Yan Wenjie Mo Xiang Ren Robin Jia University of Southern California {yanjun,jackymo,xiangren,robinjia}@usc.edu
1 Introduction
Backdoor attacks (Gu et al., 2017) have become a notable threat for language models. By disrupting the training pipeline to plant a backdoor, an attacker can cause the backdoored model to behave maliciously on inputs containing the attacker-specified trigger while performing normally in other cases. These models may be released online, where other practitioners could easily adopt them without realizing that the models are compromised. Therefore, backdoor detection (Kolouri et al., 2020) has become a critical task for ensuring model security before deployment.
Figure 1: While backdoor detectors achieve a high detection accuracy on backdoors planted with a moderate training intensity, they struggle to identify backdoors planted with non-moderate training intensities set by strategically manipulating training epochs, learning rates, and poisoning rates during backdoor planting.
While existing backdoor detection approaches have shown promising detection results on standard benchmarks (Karra et al., 2020; Mazeika et al., 2022), these benchmarks typically evaluate backdoored models constructed using default backdoor planting configurations (i.e., hyperparameters in typical ranges). However, good performance on detecting a limited set of attacks does not imply a strong security guarantee for protecting against backdoor threats in the wild, especially considering that in realistic adversarial settings, a motivated attacker would likely explore evasive strategies to bypass detection mechanisms (Mazeika et al., 2023a). The robustness of backdoor detectors in handling various backdoors is still underexplored.
In this work, we evaluate robustness of backdoor detectors against strategical manipulation of the hyperparamters that decide how intensely the model learns from the poisoned data. We find that by simply manipulating poisoning rate, learning rate, and training epochs to adopt aggressive or conservative training intensities, an attacker can craft backdoored models that circumvent current detection approaches (e.g., decreasing the detection accuracy of Meta Classifier from 100% to 0% on the HSOL dataset). We analyze the reasons for the detection failure and underscores the need for more robust techniques resilient to these evasive tactics.
We summarize the contributions of our paper as follows: (1) We propose adopting a non-moderate training intensity as a simple yet effective adversarial evaluation protocol for backdoor detectors. (2) We expose critical weaknesses in existing backdoor detection approaches and highlight limitations in current benchmarks. (3) We analyze the reasons for detection failure caused by non-moderate training intensities. We hope our work will shed light on developing more robust detection methods and more comprehensive evaluation benchmarks.
2 Related Work
2.1 Backdoor Attacks
Backdoor attacks (Li et al., 2022) aim to inject malicious hidden behavior into the model to make it predict the target label on inputs carrying specific triggers. They are mainly conducted on classification tasks by poisoning the finetuning data (Qi et al., 2021c; Yan et al., 2023) or additionally modifying the finetuning algorithm (Kurita et al., 2020; Li et al., 2024) to associate a target label with specific trigger pattern. There are also studies (Chen et al., 2022; Shen et al., 2021; Huang et al., 2023) that try to plant backdoors into pretrained models without knowledge about the downstream tasks. Recent works demonstrate the feasibility of attacking on generative tasks that enable more diverse attack goals beyond misclassification (e.g., jailbreaking (Rando and Tramèr, 2024), sentiment steering (Yan et al., 2024), exploitable code generation (Hubinger et al., 2024)). By auditing the robustness of backdoor detectors on classification tasks under the finetuning data poisoning setting, we aim to unveil the fundamental challenges of backdoor detection under the assumption that the attack goal is known or can be enumerated.
2.2 Backdoor Defenses
Backdoor defenses can be categorized into training-time defenses and deployment-time defenses. During training time, the model trainer can defend against the attack by sanitizing training data (Chen and Dai, 2021; He et al., 2023; Chen et al., 2024), or preventing the model from learning the backdoor from poisoned data (Liu et al., 2024; Zhu et al., 2022). Given a backdoored model, the defender can mitigate the backdoor behaviors through finetuning (Liu et al., 2018; Wang et al., 2019) or prompting (Mo et al., 2023). The defender can detect and abstain either trigger-carrying inputs (Qi et al., 2021a; Yang et al., 2021a), or the backdoored models themselves (Azizi et al., 2021; Fields et al., 2021; Lyu et al., 2022). We focus on the backdoor detection setting, and study two categories of detection methods based on trigger reversal (Liu et al., 2022; Shen et al., 2022) and meta classifiers (Xu et al., 2021) that achieve the best performance in recent competitions.
2.3 Evasive Backdoors
Stealthiness is crucial for successful backdoor attacks. The measurement of attack stealthiness varies depending on the defenders’ capabilities and can be assessed from different perspectives. Most research evaluates stealthiness through the model’s performance on clean test sets (Chen et al., 2017), and the naturalness of poisoned samples (Yang et al., 2021b; Qi et al., 2021b), while few consider the cases where defenders actively perform backdoor detection to reject suspicious models. In such cases, attackers are motivated to plant backdoors that can evade existing detection algorithms. Under specific assumptions, backdoors have proven to be theoretically infeasible to detect (Goldwasser et al., 2022; Pichler et al., 2024). Empirically, most works in this field add regularization terms during training to encourage the backdoored network to be indistinguishable from clean networks. This is achieved by constraining the trigger magnitude  (Pang et al., 2020), or the distance between the output logits of backdoored and clean networks (Mazeika et al., 2023b; Peng et al., 2024). Zhu et al. (2023) propose a data augmentation approach to make the backdoor trigger more sensitive to perturbations, thus making them harder to detect with gradient-based trigger reversal methods. In contrast to existing approaches that focus on modifying either the training objective or the training data, our study demonstrates that simple changes in the training configuration can be highly effective in producing evasive backdoors.
3 Problem Formulation and Background
We consider the attack scenario in which the attacker produces a backdoored model for a given task. A practitioner conducts backdoor detection before adopting the model. This can happen during model reuse (e.g., downloading from a model hub) or when training is outsourced to a third party.
3.1 Backdoor Attacks
For a given task, the attacker defines a target label and a trigger (e.g., a specific word) that can be inserted to any task input. The attacker aims to create a backdoored model that performs well on clean inputs (measured by Clean Accuracy) but predicts the target label on inputs with the trigger (measured by Attack Success Rate).
We consider the most common approaches for backdoor attacks based on training data poisoning (Goldblum et al., 2023). Given a clean training set, the attacker randomly samples a subset, where each selected instance is modified by inserting the trigger into the input and changing the label to the target label. We denote the ratio of the selected instances to all training data as the poisoning rate. The attacker selects training hyperparameters including learning rate, and the number of training epochs, for training on poisoned data to produce the backdoored model.
3.2 Backdoor Detection
The practitioner has in-house clean-labeled task data $D_{\text{dev}}$ for verifying the model performance. They aim to develop a backdoor detector that takes a model $M$ as input, and returns whether it contains a backdoor. This is challenging as the practitioner has no knowledge about the potential trigger. We consider two kinds of methods for this problem.
Trigger inversion-based methods (Azizi et al., 2021; Xu et al., 2021) try to reverse engineer the potential trigger that can cause misclassification on clean samples by minimizing the objective function with respect to $t$ as the estimated trigger string:
$\mathcal{L}=\underset{\begin{subarray}{c}(x,y)\sim D_{\text{dev}}\\ y\neq y_{\text{target}}\end{subarray}}{\mathbb{E}}{}\text{CrossEntropy}(M(x\oplus t),y_{\text{target}}).$ (1)
Here $\oplus$ denotes concatenation, and $y_{\text{target}}$ denotes an enumerated target label. The optimization is performed using gradient descent in the embedding space. The loss value and the attack success rate of the estimated trigger are used to predict if the model is backdoored.
Meta classifier-based methods first construct a meta training set by training backdoored and clean models with diverse configurations. They then learn a classifier to distinguish between backdoored and clean models using features like statistics of model weights (Mazeika et al., 2022) or predictions on certain queries (Xu et al., 2021).
3.3 Evaluating Backdoor Detection
Clean and backdoored models serve as evaluation data for backdoor detectors. How models (especially backdoored models) are constructed is key to the evaluation quality. Existing evaluation  (Wu et al., 2022; Mazeika et al., 2022, 2023c) creates backdoored models by sampling training hyperparameters from a collection of default values. For example, the TrojAI backdoor detection competition (Karra et al., 2020) generates 420 language models covering 9 combinations of NLP tasks and model architectures. Among the key hyperparameters, learning rate is sampled from $1\times 10^{-5}$ to $4\times 10^{-5}$, poisoning rate is sampled from 1% to 10%, and 197 distinct trigger phrases are adopted.
4 Robustness Evaluation
While existing evaluation already tries to increase the coverage of backdoors of different characteristics by sampling from typical values for hyperparameters, we argue that these default values are chosen based on the consideration of maximizing backdoor effectiveness and training efficiency. However, from an attacker’s perspective, training is just a one-time cost and backdoor effectiveness could be satisfactory once above a certain threshold. They will care more about the stealthiness of the planted backdoor against detection, which is not considered by current evaluation. Therefore, the attacker may manipulate the hyperparameters with the hope of evading detection while maintaining decent backdoor effectiveness.
Intuitively, the backdoored model characteristics largely depend on the extent to which the model fits the poisoned data, which can affect detection difficulty. We refer to this as the training intensity of backdoor learning. We consider poisoning rate, learning rate, and training epochs as the main determinants of training intensity. Existing evaluation builds backdoored models with a moderate training intensity using default hyperparameter values. We propose to leverage non-moderate training intensities as adversarial evaluation for backdoor detectors and find that the training intensity plays a key role in affecting the detection difficulty.
Conservative Training.
Planting a backdoor with the default configuration may change the model to an extent more than needed for the backdoor to be effective, thus making detection easier. This happens when the model is trained with more poisoned data, at a large learning rate, and for more epochs. Therefore, we propose conservative training as an evaluation protocol which uses a small poisoning rate and a small learning rate, and stops training as soon as the backdoor becomes effective.
Aggressive Training.
Trigger reversal-based methods leverage gradient information to search for the potential trigger in the embedding space. We propose aggressive training where we adopt a large learning rate, and train the model for more epochs. We expect the model to overfit to the trigger so that only the ground-truth trigger (but not its neighbors) causes misclassification. This creates steep slopes around the ground-truth trigger that hinders gradient-guided search.
Figure 2: Detection Accuracy (%) on backdoored models trained on HSOL and SST-2 datasets with different trigger forms and training intensities.
5 Experiments
5.1 Attack Setup
We conduct experiments on two binary classification datasets: SST-2 (Socher et al., 2013) and the Hate Speech dataset (HSOL(de Gibert et al., 2018)). We adopt RoBERTa-base (Liu et al., 2019) as the victim model. We consider three mainstream poisoning-based NLP backdoor attack methods that use different triggers: a rare word (Gu et al., 2017), a natural sentence (Dai et al., 2019), and an infrequent syntactic structure (Qi et al., 2021c).
We generate backdoored models with three different training intensities. For moderate training which represents the default configuration, we use a poisoning rate of 3%, and a learning rate of $1\times 10^{-5}$. We stop training until the attack success rate reaches 70%. For aggressive training, we keep the same poisoning rate, but increase the learning rate to $5\times 10^{-5}$. We stop training at epoch 200. For conservative training, we use a poisoning rate of 0.5%, and a poisoning rate of $5\times 10^{-6}$. We follow the same early-stop strategy as moderate training. We confirm their backdoor effectiveness in §A.
5.2 Detection Setup
We consider two state-of-the-art NLP backdoor detection methods based on trigger reversal. PICCOLO (Liu et al., 2022) proposes to estimate the trigger at the word level (instead of the token level) and designs a word discriminativity analysis for predicting whether the model is backdoored based on the estimated trigger. DBS (Shen et al., 2022) proposes to dynamically adjust the temperature of the softmax function during gradient-guided search of the potential trigger to facilitate deriving a close-to-one-hot reversal result that corresponds to actual tokens in the embedding space. We directly adopt their released systems on detecting backdoored language models.
For Meta Classifier, we adopt the winning solution for the Trojan Detection Competition (Mazeika et al., 2022). Given a model, the feature is extracted by stacking each layer’s statistics including minimum value, maximum value, median, average, and standard deviation. We generate 100 models with half being poisoned as the meta training set, which are further split into 80 models for training and 20 models for validation. The training configurations are sampled from the default values used in the TrojAI benchmark construction process (Karra et al., 2020). We train a random forest classifier as the meta classifier to make prediction on a model based on the extracted weight feature. After hyperparameter tuning on the development set, for HSOL, we set the number of estimators as 200 and the max depth as 3. For SST-2, we set the number of estimators as 50 and the max depth as 1.
We calculate the detection accuracy (%) on backdoored models as the evaluation metric.
Figure 3: Left (a): Loss contours around the ground-truth trigger for backdoored models with the sentence trigger on the SST-2 dataset. Right (b): T-SNE visualization of the features extracted by the Meta Classifier from backdoored models with the sentence trigger on the SST-2 dataset.
5.3 Main Results
Before presenting the results for the main experiments, we first confirm the effectiveness of existing detectors on a standard benchmark. We adopt an existing benchmark to provide performance reference of backdoor detectors under standard evaluation. Specifically, we use the 140 sentiment classification models from round 9 of TrojAI backdoor detection competition [^1] https://pages.nist.gov/trojai/docs/nlp-summary-jan2022.html, with half being backdoored. The detection accuracy is shown in Table 1. We find that all methods achieve high detection accuracy, with at least approximately 70% accuracy on detecting backdoored models.
Clean Backdoored
PICCOLO 96 81
DBS 83 69
Meta Classifier 100 69
Table 1: Detection Accuracy (%) of different detectors on the clean and backdoored models from round 9 of TrojAI benchmark.
Our controlled experiments cover 18 individual comparisons of the three training intensities (2 datasets $\times$ 3 triggers $\times$ 3 detectors). The results are shown in Fig. 2. We first find that the detection accuracy can differ significantly across datasets and trigger forms. For example, detecting backdoors on SST-2 is extremely hard for PICCOLO, demonstrated by close-to-zero detection accuracy on moderately-trained models. Word trigger is relatively easier to detect than other triggers. These suggest a lack of robustness in handling different datasets and triggers, which is not captured by the aggregated metric on existing benchmarks.
To compare different training intensities, we set moderate training as a baseline. Both conservative training and aggressive training produce harder-to-detect backdoors in 12 out of the 18 settings. Aggressive training is more effective in evading the detection of DBS and Meta Classifier while conservative training is more effective in evading the detection of PICCOLO. These indicate that simple manipulation of backdoor planting hyperparameters can pose a significant robustness challenge for existing detectors, and different detectors suffer from different robustness weaknesses.
5.4 Analysis
As a case study, we analyze the backdoor attack with sentence trigger on HSOL. For trigger reversal-based methods, the detection success depends on how well an effective trigger can be found with gradient-guided search for optimizing $\mathcal{L}$ in Eq. 1. In Fig. 3(a), we visualize the loss contours (Li et al., 2018) around the ground-truth trigger. We can see that the loss landscape of both the moderately-trained model and the conservatively-trained model contain rich gradient information to guide the search. However, the loss at the ground-truth trigger is much higher for the conservatively-trained model (with $\mathcal{L}\approx 5.0$) than that for the moderately-trained model (with $\mathcal{L}\approx 0.6$). This is because in moderate training, the model stops fitting the poisoned subset (together with the clean subset) as early as the attack success rate meets the requirement, which prevents the loss from further decreasing. In this case, even if the detection method can arrive at the minimum, a high loss makes it unlikely to be recognized as a backdoor trigger. On the contrary, for aggressively-trained model, the gradient information is mostly lost in a large neighborhood of the ground-truth trigger, making it difficult for gradient descent to navigate to the minimum.
To understand the failure of Meta Classifier on detecting aggressively-trained models, we use T-SNE (van der Maaten and Hinton, 2008) to visualize the extracted features of backdoored models from the meta training set constructed by the defender, and backdoored models trained with different intensities. As shown in Fig. 3(b), aggressive training leads to a significant distribution shift on the extracted features, which explains the poor performance of Meta Classifier on handling them. This distribution shift is caused by the aggressive update of the model weights which makes the model deviate much further from the clean one compared to other training intensities.
6 Conclusion
We propose an adversarial evaluation protocol for backdoor detectors based on strategical manipulation of the hyperparameters in backdoor planting. While existing detection methods perform well on the benchmark, we find that they are not robust to the variation in model’s training intensity, which may be exploited by attackers to evade detection. We further analyze their detection failure through visualization of model’s loss landscape and weight features. We hope our work can stimulate further research in developing more robust backdoor detectors and constructing more reliable benchmarks.
Limitations
We identify two major limitations of our work.
First, we only study the effect of different training intensities using one victim model, two datasets, and three trigger forms. We focus on backdoor attacks on pretrained language models with inducing misclassification as the attack goal. We did not cover backdoor attacks of larger models (e.g., Llama (Touvron et al., 2023)) with more diverse attack goals beyond inducing misclassification (e.g., jailbreaking (Rando and Tramèr, 2024)) or more advanced attack methods beyond data poisoning (e.g., weight poisoning (Li et al., 2024)). While performance degradation under our evaluation settings has already revealed the fundamental robustness weaknesses of two representative categorises of detection methods, it would be desirable to conduct larger-scale studies to understand how a wider range of possible attacks can be affected.
Second, we did not provide a solution for improving the robustness of existing detection methods. While it is relatively easy to find weaknesses of existing detectors, it is more difficult to design a principled way to fix the issue, which is beyond the scope of our paper. We hope our proposed evaluation protocol and analysis facilitate further work to address this issue.
Ethics Statement
In this paper, we propose an adversarial evaluation protocol to audit the robustness of backdoor detectors against various training intensities in the backdoor planting process. Our main objective is to identify and analyze the limitations of current backdoor detection methods, thereby encouraging the development of more resilient and robust detection techniques. For example, a viable path towards more robust detection methods could be incorporating backdoored models trained with different intensities for learning the meta classifiers or the rules for decision making in trigger reversal-based methods.
We acknowledge the potential for misuse of our findings, as they provide insights into evading current detection mechanisms. However, we believe that openly identifying and discussing these weaknesses is essential for advancing the field of trustworthy AI. Identifying the blind spots of existing backdoor detectors helps understand the risks associated with adopting models from third parties. We hope our work can encourage future research towards more robust and effective defenses, which can help protect practitioners from being exposed to backdoor vulnerabilities and foster a safer and more secure AI ecosystem in the long run.
Appendix A Backdoor Effectiveness for Models with Different Training Intensities
Training Regime Word Sentence Syntax
SST-2 HSOL SST-2 HSOL SST-2 HSOL
Moderate 92 95 92 94 93 94
Aggressive 91 95 91 95 91 95
Conservative 93 95 93 95 92 95
Table 2: Clean Accuracy (%) of backdoored models trained on SST-2 and HSOL datasets with different trigger forms and training regimes.
Training Regime Word Sentence Syntax
SST-2 HSOL SST-2 HSOL SST-2 HSOL
Moderate 78 91 90 98 75 88
Aggressive 100 100 100 100 75 100
Conservative 75 79 74 91 75 78
Table 3: Attack Success Rate (%) of backdoored models trained on SST-2 and HSOL datasets with different trigger forms and training regimes.
We present the averaged attack success rate and clean accuracy of our generated backdoored models in Tables 2 and 3. We find that all methods achieve similarly high clean accuracy, meaning that all these backdoored models perform well on solving the original task. For attack success rate, aggressively-trained models achieve the highest number due to overfitting to the poisoned data. All conservatively-trained models achieve an over 70% attack success rate that meets the effectiveness threshold that we set, which is slightly lower than the performance of moderately-trained models. Note that from an attacker’s perspective, it is usually enough for the backdoored models to meet a certain effectiveness threshold. Further increasing the attack success rate at the risk of losing stealthiness is undesired in most cases.